OpenStudy (anonymous):

Solve this system algebraically. 7x - 2y = 4 5y + 3x = 10 {(-41/40, -41/58)} {(41/40, 41/58)} {(-40/41, -58/41)} {(40/41, 58/41)}

OpenStudy (hba):

Did you try to solve it ?

OpenStudy (hba):

Its Just that you have to plug in the points :)

OpenStudy (anonymous):

im not sure wich one ithe right ones

OpenStudy (hba):

Well When you will plug in the values and it comes equal then you will know which one is right

OpenStudy (anonymous):

well i guess im on my own from here then

OpenStudy (anonymous):

{(-41/40, -41/58)} {(41/40, 41/58)} {(-40/41, -58/41)} {(40/41, 58/41)} are the answer choices.

OpenStudy (callisto):

For multiple choice (MC)question, putting the values into the given equations to see if it matches would be a way to do it. But knowing how to solve it is better - in case it's not a MC question. For this question, I would prefer using elimination method. 7x - 2y = 4 --(1) 5y + 3x = 10 => 3x + 5y = 10 --(2) (1) x 5 + (2) x 2 <- to get y eliminated. 5 ( 7x - 2y) + 2 ( 5y + 3x ) = 5 (4) + 2 ( 10 ) <- solve it, you can get the value of x. Then, sub the value of x into one of the given equations to get y.

OpenStudy (anonymous):

Well i guess im just gonna have tio get the problem wrong cause it gave me those as answer choices